# WildFireCan-MMD: A Multimodal dataset for Classification of User-generated Content During Wildfires in Canada

### v1

@@ 4 Classification Methodology
This section describes a range of algorithms evaluated for the 13-class classification of WildFireCan-MMD, utilizing text and image modalities. We also experimented with using Optical Character Recognition (OCR) to further extract more information from the images and create another channel of text to aid in classification. We experiment with using VLMs as zero-shot classifiers, and compare these results with fine-tuned transformer-based classifiers as well as to baseline models such as Random Forest, and Gaussian Naive Bayes. We present the implementation details of classification algorithms in the following.
4.1 VLM as Zero-Shot Classifiers
We crafted classification prompts for both gpt-4o-mini and LLaVa in an ’all-at-once’ method, meaning that with each query to the model, a list of all categories and their descriptions are given as a prompt, along with a sample (text + image) from the data to classify. We specify that the model’s task is to choose one of the provided categories and respond only with its corresponding letter. The system and user prompts can be seen in Table C.1.
4.1.1 gpt-4o-mini:
We used the available API for querying gpt-4o-mini 333version:gpt-4o-mini-2024-07-18. A specified content field in each query contains our system prompt, text from the Tweet, and user prompt. The image for the tweet is also sent with this payload.
The ’temperature’ is specified as being $0.1$ for reproducibility. ’max_tokens’ is set to be $1024$, as the text plus the image plus the model’s response will never be greater than this. The responses are saved as a JSONL file where each line contains the question ID from the questions file, as well as the model’s answer supplied as a single letter from the options shown in Table C.1. This then allowed us to calculate metrics such as balanced accuracy score and weighted precision, recall, and F1-score.
4.1.2 LLaVA:
As an open source model, querying LLaVA required a different approach than gpt-4o-mini. For our work, we opted for the 13B 4-bit version of LLaVa. The temperature was again set to $0.1$, and max_tokens set to $1024$. A custom conversation template was created using the original LLaVA codebase to construct our system prompt using the roles USER and ASSISTANT. The question file was created using the steps discussed above for gpt-4o-mini, the only difference being the addition of ’<>’ characters around each option as recommended by the source [^4] https://github.com/haotian-liu/LLaVA/blob/main/docs/Evaluation.md for querying LLaVA with a multiple-choice prompt. The text from the tweet is passed similarly to the method described above for gpt-4o-mini, and the image is converted to a tensor and passed with this to the model. The answers are saved similarly as described for gpt-4o-mini and saved, allowing us to calculate metrics such as balanced accuracy score and weighted precision, recall, and F1-score.
@@ 4.2 Custom Classifiers
To create a custom multimodal classifier for our dataset, the following steps were taken to implement concatenative early and late fusion.
@@
Early and late fusion 2-head structure::This model consists of two architectures working in parallel, an image head and a text head. The image head uses the ViT transformer, and the text head uses the RoBERTa transformer. Another linear layer is added for prediction. Figure D.1 shows the structure of our text+image ’2-head’ multimodal classifier using concatenative early fusion. We created another model with the same strategy but utilizing late fusion instead, where each head has its own final linear layer and functions as an independent classifier. The two heads output probabilities, which are then averaged for a final result.
@@
Training and Evaluation: A random seed was used for training reproducibility. Some notable hyperparameters used include a batch size of 16, an image size of ($384\times{384}\times{3}$) (as required by google/vit-base-patch16-384), and a max length for the text of 200. Twitter posts can have a maximum length of 200 characters, but after viewing the average and max length of all posts within the dataset, we learned that most users don’t go anywhere near this limit, and so this value could be decreased. We set the maximum number of epochs to be 20, and used Adam optimizer with a learning rate of $5e^{-5}$, and cross-entropy loss. Our experimentation found that the model converged at around 10-15 epochs. The model was then tested on the test set (9̃00 samples), and metrics were calculated, including balanced accuracy score and weighted precision, recall, and F1-score.
Additional Information with OCR: Trying to improve our classification metrics without gathering and labeling more data, we considered other ways in which we could extract more information from the data we had. Something we noticed in much of the data, specifically the Evacuees, Smoke & Air Quality, General Information, Preparedness, Warnings and Status Updates, and Weather Reports classes, was that many of the images contained valuable text which we believed could aid in classification. Although ViT and RoBERTA both use the transformer architecture, the way text is processed differs from images and doesn’t allow for ViT to differentiate text within images from the surrounding pixels in this current implementation.
This led us to consider ways in which we could extract text from the images in our dataset using the EasyOCR tool [^5] https://github.com/JaidedAI/EasyOCR. EasyOCR works by taking an image, preprocessing it, then applying the CRAFT algorithm to find bounding-box regions of text in an image. The pixels in these regions are sent to a text recognition model composed of a CRNN with three components: feature extraction, sequence labeling, and decoding to extract the text from these regions and turn them into strings as output. Deploying this model on our dataset, we extracted all of the text from every image and saved it to a new column as a pre-processing step. As a result, our dataset contains four columns: text, extracted_text, images, and labels. If an image was found to not have any text, a filler sentence was used: ’Image does not contain text.’.
Utilizing this extra column of text data, we added another RoBERTa head to our classifier. This new classifier, now with three separate heads concatenated and fed into a single linear layer, as appeared in Figure D.2. Note that this figure represents the early fusion version. A late fusion version was also created similarly as described for the 2-head classifier. The model was then trained with the same settings as described previously, for a similar number of epochs.
@@ 5 Experimental Results
Table 2 shows the calculated F1-scores for the VLMs, custom classifiers, and two baseline models per class and as an average over all classes. Also, figure E.1 shows the confusion matrices of the VLMs and the custom classifiers. Note that all metrics reported in these tables use ’weighted’ F1-scoring to better represent the class imbalance observed in our dataset.
Table 2: F1-scores per-class for VLMs, custom classifiers, and baseline models.
@@
Class LLaVa gpt-4o-mini 2-head-early 3-head-early 2-head-late 3-head-late R. Forest Gaussian NB
Evacuees 0.17 0.37 0.77 0.73 0.79 0.74 0.29 0.50
General Information 0.14 0.16 0.59 0.46 0.53 0.59 0.35 0.67
Preparedness 0.61 0.65 0.82 0.81 0.86 0.84 0.52 0.80
Weather Reports 0.59 0.64 0.88 0.91 0.91 0.91 0.45 0.66
Warnings & Status Updates 0.36 0.45 0.80 0.76 0.77 0.74 0.50 0.57
Reports of Actions of Responders 0.29 0.63 0.89 0.86 0.85 0.83 0.48 0.74
Infrastructure 0.38 0.40 0.92 0.95 0.90 0.93 0.62 0.64
Political 0.70 0.77 0.84 0.84 0.84 0.85 0.28 0.78
Insurance 0.93 0.90 0.93 0.89 0.93 0.91 0.92 0.91
Advertisement 0.08 0.68 0.76 0.76 0.74 0.75 0.23 0.81
Smoke and Air Quality 0.76 0.79 0.86 0.86 0.86 0.85 0.58 0.70
Support 0.25 0.55 0.73 0.63 0.75 0.78 0.45 0.63
Other 0.05 0.54 0.78 0.78 0.76 0.79 0.36 0.60
Average 0.41 0.58 0.81 0.79 0.81 0.81 0.46 0.69
Weighted Average 0.45 0.60 0.83 0.81 0.82 0.82 0.48 0.67
LLaVa has limited capabilities.Testing on WildFireCan-MMD, We observed that LLaVa is not able to reach an acceptable result, performing close to the Random Forest baseline and underperforming gpt-4o-mini by a large margin. Specifically, we noted a large difference in F1-score between gpt-4o-mini and LLaVA for the Advertisement and Other classes (gpt4o-mini achieved between 0.5 and 0.6, whereas LLaVA reaches only less than 0.1). This can be explained by the fact that these two classes have the highest variance by far and are hard to handle for LLaVa. Specifically, the Other category is very diverse since tweets are labeled as so if they don’t fit into any of the other categories. Referring to confusion matrices (Figure E.1(a)), we can see that LLaVA struggles with labeling the Warnings & Status Updates class, where it often classified tweets from other classes as such, especially the Other class.
gpt-4o-mini shows strengths and limitations across various categories. On the other hand, gpt-4o-mini, being much larger and more sophisticated than LLaVA, outperforms it in almost every class except for the Insurance category, where it has been beaten by LLaVA by a couple of percent. Interestingly, gpt-4o-mini reaches a similar F1-score as custom classifiers on the Smoke & Air Quality, Advertisement, Insurance and Political classes, which combined encompass over a third of the dataset. This is probably due to the lower variance of these classes. On the contrary, the classes in which it suffers the most are Evacuees, General Information, Warnings & Status Updates, Infrastructure which contain some of the highest variance. It is difficult to construct a prompt, which encompasses the variance of these classes, while staying within the token limit. So, this may be one reason why gpt4-o performed worse than custom models in these categories. As seen in Figure E.1(b) we can see that gpt-4o-mini had the most trouble with the Warnings & Status Updates class, where it often confused it with the General Information and Evacuees classes. This effect caused poor performance in the Infrastructure class as well.
Custom classifiers are more robust and reliable. Referring to Table 2, we can see that the two and three-head classifiers provide similar results. Our 2-head classifiers have an average weighted F1-score of 0.83 for early fusion, and 0.82 for late fusion. The 3-head classifiers have an average weighted F1-score of 0.81 for early fusion, and 0.82 for late fusion. Surprisingly, the three-head classifier performs slightly worse in most classes than the two-head classifier, even though it receives extra contextual information from the images. One important consideration, however, when thinking about the utility and purpose of this work, is that the 3-head classifiers perform slightly better on key classes such as Infrastructure, Weather Reports, Smoke & Air Quality, General Information, and Support. Infrastructure is an important class because it can provide information to first responders and administrative personnel about key infrastructure that has been damaged or is currently in need of focus. Weather Reports is also important because it can provide insights and predictions about weather patterns and likely areas of new fires. General Information is useful to administrators and the general public as this class contains broad information about the situation as a whole.
@@ 6 Discussion
The use of VLMs involves tradeoffs. Practical implications are involved in benchmarking with VLMs in a zero-shot setting. We discuss these implications from three perspectives: accuracy, cost and compliance with the responsible AI principles. The closed-source model, gpt-4o-mini, while being accurate to some extent, is expensive for large-scale deployment. Also, closed-source models often fail to comply with responsible AI principles due to a lack of transparency, privacy, and limited scrutiny and ethical oversight Hua et al. (2024). On the other hand, despite the free and open-source nature of LLaVA, its performance falls short in critical tasks such as disaster management. If performance is improved, the open-source nature of this model might make it suitable for disaster response, allowing compliance with responsible AI guidelines through offering access to the code base and ensuring ethical scrutiny. Particularly, the transparent nature of open source is crucial when handling sensitive data that may include personal information. Despite gpt-4o-mini performing better on our test set, it is important to keep these above considerations in mind. When labeled data is unavailable, VLMs can function as a zero-shot classifier, but in practice, neither model is particularly well suited to this specific type of task, emphasizing the necessity for a custom fine-tuned model.
Fine-tuning on task-specific data is essential. Results presented in Table 2 demonstrate that even simple approaches such as Random Forest and Gaussian NB, when trained with task-specific labelled data, provide similar or better results than what is achieved from zero-shot VLMs. Even though our dataset is modest in size, it clearly showcases the benefits of developing a labeled dataset and fine-tuning a custom classifier. Further, our results show that high degrees of accuracy can be achieved for this task by fine-tuning advanced classification structures, such as pre-trained transformer-based text and image encoders.
The custom classifiers we have developed serve as a foundational step for triaging tweets during wildfires, providing an initial performance benchmark and demonstrating our approach’s feasibility. However, to enhance the accuracy and robustness of our model, future work will need to adopt a more comprehensive approach to experimentation. One area for improvement is the application of different fusion strategies, such as multiplicative fusion, which has been shown to better capture the relationship between different modalities Koshy and Elango (2023).
@@ 7 Conclusion
General-purpose VLMs, while powerful, struggle to perform reliably in high-stakes, context-rich scenarios like disaster response. To address this gap, we present WildFireCan-MMD, a wildfire-specific, multimodal dataset of social media posts (text + image) curated from X, accompanied by a domain-focused taxonomy derived through multimodal topic modeling and manual annotation. Unlike prior disaster datasets, ours is tailored exclusively to wildfire contexts, making it a valuable benchmark for evaluating true generalization in VLMs. As climate-induced wildfires grow in frequency and complexity, this dataset supports the development of robust, transparent, and adaptive AI tools for detection, coordination, and recovery.

### v2

@@ 4 Classification Methodology
This section describes a range of algorithms evaluated for the 13-class (12 categories of the taxonomy +“Other”) classification of WildFireCan-MMD. In general, we employ VLMs as zero-shot classifiers, and compare the results with fine-tuned transformer-based classifiers as well as to classical methods such as decision tree, and Gaussian naive bayes, as minimum baselines.
Since each post on X can have up to 4 associated images, we duplicated the text and paired with each image as an individual sample. For training and testing the custom classifiers, we split the data into an 80/20 train/test split using stratified sampling, to ensure similar distributions of classes in train and test subsets. We use the same test set for all of the classifiers, including zero-shot VLMs.
4.1 VLMs as Zero-Shot Classifiers:
We considered three VLMs, GPT-4 Achiam et al. (2023), LLaVA Liu et al. (2023), and SmolVLM Marafioti et al. (2025) and tested them in a zero-shot setting. We created classification prompts for all three VLMs where, with each query to the model, a list of all categories and their descriptions is given as a prompt along with a data instance (text + image) to classify. We specify that the model’s task is to choose one of the provided categories and respond only with its corresponding letter. The system and user prompts can be seen in Table C.1.
gpt-4o-mini: We used the available API for querying gpt-4o-mini 333version:gpt-4o-mini-2024-07-18. A specified content field in each query contains our system prompt, text and image from the tweet, and user prompt. The responses are saved as a file where each line contains the question ID from the questions file, as well as the model’s answer supplied as a single letter from the options shown in Table C.1. This then allows for the calculation of metrics.
LLaVA: We opted for the 13B 4-bit version of LLaVA due to hardware constraints. Using a combination of a publicly available Colab notebook  [^4] https://github.com/camenduru/LLaVA-colab and the original code from the open-source LLaVA repository  [^5] https://github.com/haotian-liu/LLaVA, the model was loaded onto the GPU within Colab. The answers are saved similarly as described for gpt-4o-mini.
smolVLM: Being the smallest of the VLMs experimented with, only requiring <10GB of GPU ram, experiments were able to be performed locally without quantization. Due to its inclusion in the Transformers library, loading and querying the model with the previously created prompts was simpler and cheaper than the other methods, while taking a similar amount of time to process.
@@ 4.2 Custom Classifiers
Prior work by Koshy and Elango (2023) demonstrated that a multimodal architecture combining RoBERTa with a BiLSTM and ViT, using multiplicative fusion and an additional attention mechanism, achieved the best performance on the CrisisMMD dataset. We evaluated this architecture, among several others, on our WildFireCA-MMD dataset. Our findings indicate that the BiLSTM and additional attention components did not improve performance in our context. Instead, a simpler architecture employing early fusion via feature concatenation of RoBERTa and ViT embeddings yielded the best results. We experimented with several model structures described in the following.
@@
2-head structure:This model consists of two parallel architectures, an image head and a text head. The image head uses the ViT666’google/vit-base-patch16-384’ transformer, and the text head uses the RoBERTa777’roberta-base’ transformer. A linear layer is added for prediction. Figure D.1 shows the structure of our text+image ’2-head’ multimodal classifier using concatenative early fusion.
3-head structure: To enhance classification without more manual labeling, we extracted text from images using EasyOCR [^8] https://github.com/JaidedAI/EasyOCR. The extracted text was saved in a new column (extracted_text), defaulting to a filler if none was found. Our classifier was updated to include a third RoBERTa head for this column, as shown in Figure D.2.
@@
Training and Evaluation: A random seed was used for reproducibility. Some notable hyperparameters used include a batch size of 16, an image size of ($384\times{384}\times{3}$) (as required by our ViT model). We set the maximum number of epochs to be 20, and used Adam optimizer with a learning rate of $5e^{-5}$, and cross-entropy loss. Our experimentation found that the model converged at around 10-15 epochs. The model was then tested on the test set (~1000 samples), and metrics were calculated, including balanced accuracy score and weighted precision, recall, and F1-score.
4.3 Baseline Classifiers
To establish baseline, we evaluated several classic classifiers on our dataset, namely Decision Tree, Gaussian Naive Bayes, and K-Nearest Neighbors (KNN). For textual data, we extracted TF-IDF feature matrices by applying the TfidfVectorizer with max_features=5000. The image data, was resized to 64$\times$64 pixels. The resulting TF-IDF and image matrices were concatenated and used as input to the classifiers. Hyperparameters were set to default values in Scikit-learn  [^9] https://scikit-learn.org/stable/.
@@ 5 Experimental Results
Table 2 shows the calculated F1-scores for the VLMs, custom classifiers, and baseline models per class and as an average over all classes. Also, Figures E.1 and E.2 show the confusion matrices of the custom classifiers and VLMs, respectively. Note that all metrics reported in these tables use ’weighted’ F1-scoring to better represent the class imbalance observed in our dataset.
Some classes are harder to distinguish than others. In general, we observe a large variance of classification F-scores across classes. For example, all classifiers struggle with classifying examples of Evacuees and General Information. We explain this observation by the great variability of information in these classes. On the other hand, classes with less information variety, such as Insurance and Smoke and Air Quality, are more manageable for most classifiers to detect.
Table 2: F1-scores per-class for VLMs, custom and baseline classifiers. Maximum results are in bold.
@@
Class LLaVA GPT-4o-mini smolVLM 2-head 3-head Dec. Tree Gauss. NB KNN
Evacuees 0.17 0.37 0.25 0.77 0.73 0.47 0.53 0.29
General Information 0.14 0.16 0.00 0.59 0.46 0.23 0.46 0.21
Preparedness 0.61 0.65 0.23 0.82 0.81 0.57 0.72 0.49
Weather Reports 0.59 0.64 0.00 0.88 0.91 0.46 0.63 0.52
Warnings & Status Updates 0.36 0.45 0.00 0.80 0.76 0.45 0.59 0.43
Reports of Actions of Responders 0.29 0.63 0.00 0.89 0.86 0.45 0.70 0.40
Infrastructure 0.38 0.40 0.00 0.92 0.95 0.47 0.75 0.52
Political 0.70 0.77 0.00 0.84 0.84 0.52 0.75 0.31
Insurance 0.93 0.90 0.92 0.93 0.89 0.93 0.93 0.89
Advertisement 0.08 0.68 0.00 0.76 0.76 0.22 0.65 0.40
Smoke and Air Quality 0.76 0.79 0.50 0.86 0.86 0.60 0.69 0.51
Support 0.25 0.55 0.06 0.73 0.63 0.30 0.60 0.33
Other 0.05 0.54 0.27 0.78 0.78 0.35 0.67 0.33
Weighted Average 0.45 0.60 0.20 0.83 0.81 0.48 0.67 0.43
VLMs struggle with nuanced disaster communication, underperforming compared to classic classifiers. Table 2 shows that VLMs, including LLaVA, GPT-4o-mini, and smolVLM, consistently underperform across most disaster-related classes. Their weighted F1-scores range from 0.20 (smolVLM) to 0.60 (GPT-4o-mini), all falling short of the best-performing baseline (Gaussian Naive Bayes, 0.67). Similarly, the Decision Tree classifier achieved a weighted score of 0.48, comparable to LLaVA’s 0.45. The performance of smolVLM in particular was notably poor, with a weighted average of only 0.20 and zero scores in multiple categories. These results suggest that, despite their general-purpose capabilities, current VLMs struggle to perform reliably in domain-specific, fine-grained classification tasks, such as the case for WildFireCan-MMD.
Custom classifiers are more robust and reliable. In contrast to VLMs, our custom classifiers (2-head and 3-head architectures) demonstrate superior and more stable performance. With weighted F1-scores of 0.83 and 0.81, they outperform all baselines and VLMs across nearly all categories. The 2-head classifier achieves top scores in Insurance (0.93), Reports of Actions of Responders (0.89), Smoke and Air Quality (0.86). The 3-head model benefits from incorporating OCR-extracted text, performing particularly well in Infrastructure (0.95), and Weather reports (0.91). While incorporating OCR-extracted textual information doesn’t improve the classification results overall, it might be a favorable approach when certain classes are deemed to be more critical for certain response strategies.
6 Analysis of Trends in Data
After labeling our dataset and determining our best-performing classifier (the 2-head classifier described in section 4.3 and Figure D.1), we proceeded to collect another 46,279 posts from X. The collection methodology was the same as explained in section 3.1, where we gathered tweets using a curated set of wildfire-related hashtags specific to each year, selected based on known wildfire events and hotspot locations across affected regions. We combined this with a keyword-based search, then combined the results and removed duplicates. The full set of queries used for collecting this new data can be seen in Appendix F.
First, we extracted the author locations from the samples. Our goal was to obtain the province-level region where each sample is coming from. After extracting the author locations, which could be either automatically generated or user-provided, we ran the queries through OSMNX [^10] https://osmnx.readthedocs.io/en/stable/. This allowed us to automatically search for the locations of cities or towns that users provided and extract the province level information. If a users location was not found, or not in Canada, it was labeled as such. The distribution of posts by province can be seen in Figure 2, showing that the majority of data is posted from British Columbia and Alberta. Additional information about this data, including a word cloud of the text and the frequency distribution of the posts by word count can be found in Figures 1(a) and 1(b).
Figure 2: Distribution of collected tweets based on author location.
Using our best-performing classifier with an F1-score of 83%, we obtained predictions for the new data. This automated classification approach allowed for efficiency in processing the large volume of newly collected data without having to manually label each individual sample. This of course, will not uphold the same gold standard of labeling that a human annotator could achieve, and manual labeling will have to be done in the future. Individual classifications should be approached with skepticism, but this method should still allow for general trends in the data to be analyzed on a broader scale.
Figure 3 (left) shows the total number of posts per week sorted by year. This demonstrates a significant event occurred the week of June 3, 2023, and July 22, 2024, due to several major fires in the country Reuters Staff (2023); CBC News (2024). Figure 3 (right) illustrates weekly trends in predicted labels during 2023. Sharp spikes in Smoke and Air Quality and Warnings & Status Updates likely correspond to major wildfire events occurring in Alberta and BC. Warnings & Status Updates and Evacuees follow similar patterns around the week of Aug 12, which corresponds to the state of emergency declared as the city of West Kelowna was evacuated. Overall, at least 35,000 people were under evacuation order and another 30,000 under evacuation alert, as of August 19 Williams (2023). Reports of Actions of Responders and Support also rise during these peaks, reflecting coordinated emergency efforts.
Figure 3: Weekly post distributions: (left) overall yearly trend, (right) class-wise distribution in 2023.
@@ 7 Discussion
The use of VLMs involves tradeoffs. Practical implications are involved in benchmarking with VLMs in a zero-shot setting. We discuss these implications from three perspectives: accuracy, cost and compliance with the responsible AI principles. The closed-source model, gpt-4o-mini, while being accurate to some extent, is expensive for large-scale deployment. Also, closed-source models often fail to comply with responsible AI principles due to a lack of transparency, privacy, and limited scrutiny and ethical oversight Hua et al. (2024). On the other hand, despite the free and open-source nature of LLaVA, its performance falls short in critical tasks such as disaster management. Also, smolVLM, while being both open-source and computationally affordable, has a performance far below acceptable rates. Therefore, when labeled data is unavailable, VLMs might be considered as a zero-shot classifier, but in practice, these models are not particularly well suited to this specific type of task, emphasizing the necessity for a custom fine-tuned model.
Fine-tuning on task-specific data is essential. Results presented in Table 2 demonstrate that even simple approaches such as decision tree and Gaussian NB, when trained with task-specific labeled data, provide similar or better results than what is achieved from zero-shot VLMs. Even though our dataset is modest in size, it clearly showcases the benefits of developing a labeled dataset and fine-tuning a custom classifier. Further, our results show that high degrees of accuracy can be achieved for this task by fine-tuning common classification structures, such as pre-trained transformer-based text and image encoders. These findings align with other works such as Nadeem et al. (2025).
Custom classifiers can be used to discover social media use patterns. The primary function of the custom classifier is to perform real-time labeling of social media data during wildfire events, enabling the timely extraction of actionable information to support disaster response efforts. Beyond real-time deployment, our post-hoc analysis of a large unlabeled dataset illustrates the classifier’s capacity to identify latent patterns and temporal trends, providing a scalable tool for retrospective analysis and evidence-based policy development.
To enhance the accuracy and robustness of our model, future work can consider the application of different fusion strategies, such as multiplicative fusion, which has been shown to better capture the relationship between different modalities Koshy and Elango (2023). Fine-tuning of VLMs is also a promising avenue of improving the task which should be explored in future works Duan et al. (2024); Zhai et al. (2024).
@@ 8 Conclusion
Despite the advancements in general-purpose VLMs, they struggle to perform reliably in high-stakes, context-rich scenarios like disaster response. This low accuracy, alongside the resource-intensive nature of such models, necessitates developing smaller fine-tuned models, specialized for specific regions and contexts. To address this gap, we present WildFireCan-MMD, a wildfire-specific, multimodal dataset of social media posts curated from X, accompanied by a domain-focused taxonomy derived through multimodal topic modeling and manual annotation. Unlike prior disaster datasets, ours is tailored exclusively to wildfire contexts, making it a valuable benchmark for evaluating true generalization in VLMs, and suitable for training specialized models. As climate-induced wildfires grow in frequency and complexity, this dataset supports the development of robust, transparent, and adaptive AI tools for detection, coordination, and recovery.
9 Acknowledgment
This work was funded by the National Program Office of the National Research Council Canada, under the New Beginning Initiative Program.